3 meter long piece of wood, cut into 2 parts, 1 part is 3times longer than the 2nd part. How long is the short piece?
Answer by Cromlix(4381)

Ratio.
3:1
Add 3 + 1 = 4
1/4 * 3 = 0.75m
3/4 * 3 = 2.25m
Short piece measures 0.75m.
Hope this helps.
:-)